Exercise · ZW-BINOMIAL-DISTRIBUTION-006 · ★★☆ Intermediate
A fair coin is tossed 3 times. What is the probability of getting at least one head?
Choose one option.
Show answer
- C. \(\tfrac78\)
Exercise · ZW-BINOMIAL-DISTRIBUTION-006 · ★★☆ Intermediate
Choose one option.
Use the complement: \(1-P(X=0)=1-(1/2)^3=7/8\). The value \(1/8\) is the probability of no heads.
A binomial variable \(X\) counts the successes \(k\) in \(n\) independent Bernoulli trials with the same success probability \(p\); \(P(X=k)\) is the probability of getting \(k\) successes.
← ZW-BINOMIAL-DISTRIBUTION-005 ZW-BINOMIAL-DISTRIBUTION-007 →